OpenStudy (anonymous):

please help me with Q1.....give formula to solve

OpenStudy (anonymous):

what is r?

OpenStudy (anonymous):

1: Vol = (Vs*Ron)/(Ron+r) 2: Tr = -Rl*Cgs*ln((Vs-Voh)/Vs) 3: Rl||Ron 4: Tau = Rth*Cgs, where Rth = Ron||Rl 5: Tf = -Rth*Cgs*ln((Vol-Vth)/(Vs-Vth))

OpenStudy (anonymous):

r ????

OpenStudy (anonymous):

where is r?

OpenStudy (anonymous):

in 1: Vol=(vs*Ron)/(Ron+r)

OpenStudy (anonymous):

r=x )

OpenStudy (anonymous):

i cant get u

OpenStudy (anonymous):

1=(2*3)/(3+x) x=?

OpenStudy (anonymous):

how do u get this?

OpenStudy (anonymous):

The static discipine of the inverter is VIL=1.5V, VOL=1V, VIH=3.5V, and VOH=4V. Vol is a known value

OpenStudy (anonymous):

plz give me ans for q1

OpenStudy (anonymous):

Rl is R ?

OpenStudy (anonymous):

Rl is R

OpenStudy (anonymous):

how to find Vth?